### v1

# Breaking the Script Barrier in Multilingual Pre-Trained Language Models with Transliteration-Based Post-Training Alignment

Abstract
Multilingual pre-trained models (mPLMs) have shown impressive performance on cross-lingual transfer tasks. However, the transfer performance is often hindered when a low-resource target language is written in a different script than the high-resource source language, even though the two languages may be related or share parts of their vocabularies. Inspired by recent work that uses transliteration to address this problem, our paper proposes a transliteration-based post-pretraining alignment (PPA) method aiming to improve the cross-lingual alignment between languages using diverse scripts. We select two areal language groups, Mediterranean-Amharic-Farsi and South+East Asian Languages, wherein the languages are mutually influenced but use different scripts. We apply our method to these language groups and conduct extensive experiments on a spectrum of downstream tasks. The results show that after PPA, models consistently outperform the original model (up to 50% for some tasks) in English-centric transfer. In addition, when we use languages other than English as sources in transfer, our method obtains even larger improvements. We will make our code and models publicly available at https://github.com/cisnlp/Transliteration-PPA.
Breaking the Script Barrier in Multilingual Pre-Trained Language Models with Transliteration-Based Post-Training Alignment
**footnotetext: Equal contribution.
1 Introduction
Recent mPLMs such as mBERT (Devlin et al., 2019) and XLM-R (Conneau et al., 2020) have shown remarkable performance on cross-lingual transfer tasks by learning cross-lingual representations from monolingual corpora (Pires et al., 2019; Artetxe et al., 2020a). Despite their impressive performance, these models still exhibit limitations in cross-lingual transfer involving low-resource languages. Deshpande et al. (2022) showed that the downstream performance of mPLMs is correlated with the degree of alignment between word embeddings across languages. Another factor that hinders the knowledge transfer is the script diversity or script barrier of represented languages, which has been observed even in the case of related languages (Anastasopoulos and Neubig, 2019; Muller et al., 2021). The script barrier problem can also be viewed from the perspective of representation alignment. Wen-Yi and Mimno (2023) showed that token representations from different scripts could be almost perfectly linearly separated, indicating that models struggle to learn a common representation space. Therefore, post-training is required to boost zero-shot cross-lingual transfer in tasks like sentence retrieval, text classification, or sequence labeling, all of which benefit from better cross-lingual alignment (Hämmerl et al., 2024).
Many post-training alignment strategies use objectives that rely on bilingual dictionaries or parallel data to align the representations of mPLMs (Cao et al., 2020; Wang et al., 2020; Schuster et al., 2019; Pan et al., 2021). However, dictionaries and parallel corpora are often limited in the data scale or the number of languages they cover (Artetxe et al., 2020b), which might be impractical for building strong supervision signals for many low-resource languages. Another alternative that improves cross-lingual alignment is to use transliteration (a process of converting the text of a language from one script to another). Transliteration can improve the lexical overlap, especially for related languages (Moosa et al., 2023). Different from translation, transliterations can be obtained nearly for free using well-performing rule-based tools (Hermjakob et al., 2018). Therefore, several works have shown improvements in cross-lingual transfer by pre-training or fine-tuning models with data transliterated into a common script (Murikinati et al., 2020; Muller et al., 2021; Purkayastha et al., 2023; Moosa et al., 2023). However, they require the use of a single common script by the model. This is restrictive for many tasks as the process of transliteration can be lossy and non-invertible.
Recently, Liu et al. (2024a) proposed a sequence-level contrastive learning objective to improve the alignment across different scripts at a large scale (for more than 500 languages), using sentences in both their original scripts and their Latin-script transliterations, validated by English-centric cross-lingual zero-shot transfer evaluations. However, there are three major limitations in their setup. First, the contrastive objective only manipulates the sequence-level representations in the middle layer, which does not directly contribute to better alignment in the token-level space. Second, not every language pair has extensive lexical overlap that can boost cross-lingual transfer: transliteration-based alignment makes more sense for mutually influenced languages. Lastly, English alone as a transfer source language does not fully exploit the alignment benefit, as it does not have the most lexical overlap with other languages.
To this end, we propose a new transliteration-based post-training alignment method that works on both sequence and token levels. Our method does not rely on parallel data. Instead, similar to Liu et al. (2024a), we use the monolingual data in their original scripts and their Latin transliteration obtained by using Uroman (Hermjakob et al., 2018), a rule-based transliteration tool. We investigate the impact of the strategy by focusing on two groups of languages: Mediterranean-Amharic-Farsi and South+East Asian Languages, described more in detail in Section 4.2. The languages in each group share areal features but differ in scripts. Some languages are closely related as members of the same language family (e.g., Semitic and Sino-Tibetan). Additionally, languages in each group have extensive lexical overlap due to historical contact and geographical proximity (e.g., Chinese and Korean or Turkish and Arabic). As these languages are written in different scripts, transliteration can help to better exploit the shared linguistic properties and thus improve the cross-lingual transfer.
We leverage our method to post-train Glot500 (ImaniGooghari et al., 2023) (a continually pre-trained model from XLM-R on more than 500 languages) on the selected language groups and evaluate the zero-shot cross-lingual transfer performance on three types of downstream tasks: sentence retrieval, text classification, and sequence labeling. The evaluation is done with English as the source language, as well as with three other source languages of different scripts in each group. We show that our method consistently improves the downstream task performance across different languages and scripts. Moreover, our method further boosts the transfer performance when better source languages are chosen, as the performance depends on the degree of alignment between the source and target languages – precisely the alignment that our approach boosts.
Our contributions can be summarized as follows: (i) We propose a transliteration-based post-training alignment method that operates on both sequence and token levels, aiming to bridge the script barrier in mPLMs; (ii) We investigate the impact of our method on two areal groups of languages with different scripts and show consistent improvements in zero-shot cross-lingual transfer; (iii) We systematically explore how different source languages influence the zero-shot transfer performance of our obtained transliteration-aligned models.
2 Related Work
Many recent works have proposed pre-training or fine-tuning alignment methods to improve cross-lingual transfer in mPLMs. Cao et al. (2020) proposed a fine-tuning embedding alignment objective between word pairs procured in an unsupervised fashion from parallel data using statistical word alignment models (Dyer et al., 2013). Chaudhary et al. (2020) improved alignment during pre-training by using bilingual dictionaries to replace words in original sentences with translations in other languages. Similarly, Tang et al. (2022) used bilingual dictionaries to explicitly align the embeddings of the same words in different languages during pre-training. Wei et al. (2021) proposed a hierarchical contrastive learning pre-training method, which uses parallel data to align representations at the word and sentence levels. Similarly, Hu et al. (2021) proposed a pre-training method with explicit alignment signals from parallel data that encourages symmetry at both word and sentence levels. Pan et al. (2021) combined contrastive learning with translation language modeling (Conneau and Lample, 2019) as a post-training alignment method that uses parallel data as well. While these methods have shown improvements in cross-lingual transfer, they have the limitation of requiring parallel data or bilingual dictionaries, which may be hard to acquire for many low-resource languages.
Transliteration is a process of converting the text of a language from one script to another (Wellisch et al., 1978). This process does not involve translating meanings but rather represents the original symbols as closely as possible in the target script. Different works have proposed transliteration-based methods to address the script barrier problem in multilingual models. Murikinati et al. (2020) used transliteration to a common script to improve cross-lingual morphological inflection. Khemchandani et al. (2021) exploited language relatedness between Indian languages and leveraged transliteration to a common script to adapt multilingual models to low-resource languages. Muller et al. (2021) analyzed the behavior of multilingual models on unseen languages and found that languages written in different scripts do not benefit from transfer learning. They proposed transliteration to the high-resource source language script as a solution to address the script barrier. Purkayastha et al. (2023) showed that fine-tuning multilingual models on data transliterated into Latin script improves cross-lingual transfer for low-resource languages. Similarly, Moosa et al. (2023) pre-trained models from scratch on data transliterated into a common script for the Indic languages and showed improvements in cross-lingual transfer. Our work is most related to TransliCo proposed by Liu et al. (2024a), where a sequence-level contrastive learning objective is used to encourage alignment across different scripts without restricting the models to a common script, using data in their original scripts and their Latin-script transliteration. However, their English-centric evaluation setup limits the ability to fully reveal the impact of transliteration. This paper systematically explores how transliteration-based alignment enhances the performance of using various source languages.
3 Methods
We present a transliteration-based post-training alignment method that can be used to fine-tune existing encoder-only mPLMs for improved alignment across languages using different scripts, boosting cross-lingual transfer performance. Our method consists of three objectives: masked language modeling, sentence-level alignment, and token-level alignment. All objectives are trained on combined original and transliterated data. The transliterated data is obtained by converting the original data into Latin script. The transliteration process uses Uroman (Hermjakob et al., 2018), a rule-based system that can convert nearly all character sets into a common Latin script. The overall method is illustrated in Figure 1, and we introduce the three objectives in detail in the following.
Figure 1: Overview of our transliteration-based post-training alignment method consisting of three objectives: masked language modeling ($\mathcal{L}_{\text{MLM}}^{orig}$ and $\mathcal{L}_{\text{MLM}}^{latn}$), sentence-level alignment ($\mathcal{L}_{\text{SEQ}}$), and token-level alignment ($\mathcal{L}_{\text{TLM}}$).
3.1 Masked Language Modeling
Given an input sequence in its original script: $X^{orig}_{i}$ or its transliterated version: $X^{latn}_{i}$, we apply the naive MLM objective (Devlin et al., 2019) to predict randomly masked tokens in both sequences:
$\mathcal{L}_{\text{MLM}}=\mathbb{E}\left[-\sum_{m\in\mathcal{M}}\log p_{\text{MLM}}(X_{i,m}|\boldsymbol{h}_{i,m})\right]$
where $\mathcal{M}$ is the set of masked positions in the input sentence $X_{i}$ (either $X^{orig}_{i}$ or $X^{latn}_{i}$) and $P_{\text{MLM}}(X_{i,m}|\boldsymbol{h}_{i,m})$ is the probability of predicting token $X_{i,m}$ giving $\boldsymbol{h}_{i,m}$, the final contextualized representation at the position $m$ in the $i$th sequence. The probability is computed by an MLM head. Fine-tuning with MLM on original data is necessary to preserve the model’s knowledge. On the other hand, the mPLM has very limited knowledge about the transliterated data, which makes the MLM objective on transliterated data crucial for learning useful cross-script representations. We refer to the MLM objective for the original data (resp. transliterated data) as $\mathcal{L}_{\text{MLM}}^{orig}$ (resp. $\mathcal{L}_{\text{MLM}}^{latn}$).
3.2 Sentence-Level Alignment
We treat an input sequence in the original script $X^{orig}_{i}$ and its transliterated version $X^{latn}_{i}$ as having the same semantics. Therefore, we apply a sequence-level contrastive learning objective, similar to SimCSE (Gao et al., 2021), to encourage the model to learn similar sequence-level representations for the original and transliterated sequences. This setting is analogous to other works that apply contrastive learning on pairs formed by an original sentence and its English translation (Chi et al., 2021; Pan et al., 2021). In our context, Latin acts as a pivot script, encouraging better cross-lingual alignment of representations in different scripts.
Following Liu et al. (2024a), we apply the contrastive learning objective on a given batch of original and transliterated sequences $B=\{(X^{orig}_{i},X^{latn}_{i})\}_{i=1}^{N}$. Each batch defines positive contrastive pairs $(X,X^{+})$ where $X$ is the original sequence and $X^{+}$ is its transliterated version or vice versa, i.e., $(X^{orig}_{i},X^{latn}_{i})$ or $(X^{latn}_{i},X^{orig}_{i})$. For each positive pair, the negative examples are formed by all other sequences in the batch $B^{-}=B\setminus\{(X^{orig}_{i},X^{latn}_{i})\}$ (slightly abusing notation). The contrastive loss is then defined as:
$\mathcal{L}_{\text{SEQ}}=\mathbb{E}\left[-\log\frac{e^{\text{sim}(f(X),f(X^{+}))/\tau}}{e^{\text{sim}(f(X),f(X^{+}))/\tau}+\text{NEG}}\right]$
where $\text{NEG}=\sum_{(X,X^{-})\in B^{-}}e^{\text{sim}(f(X),f(X^{-}))/\tau}$, $f$ is defined as mean pooling over the 8th layer output contextualized embeddings (ignoring the special tokens output except for [mask] token), sim is the dot product, and $\tau$ is the temperature set to $1$.
3.3 Token-Level Alignment
The sentence-level alignment objective helps the model to learn similar sentence representations for the original and transliterated sequences. This is useful for improving performance in sentence-level downstream tasks like sentence retrieval or classification. However, this objective manipulates the output of a middle layer, which does not directly contribute to better alignment in the token-level space. For token-level tasks like NER and POS tagging, alignment at the token level might be more beneficial. Therefore, we propose a token-level alignment objective that further encourages the model to align the representations of the original and transliterated words. We adapt the translation language modeling objective introduced by  Conneau and Lample (2019), which is equivalent to applying the MLM objective on a concatenated bilingual sentence pair. Specifically, given a sentence pair $(X^{orig}_{i},X^{latn}_{i})$, we apply the MLM objective on the concatenated sequence $X^{orig}_{i}\oplus X^{latn}_{i}$ or $X^{latn}_{i}\oplus X^{orig}_{i}$, where concatenation order is randomly chosen during training. The intuition is that, to predict a token masked in the original sentence, the model can either attend to surrounding tokens in the original script or their transliterations and vice versa. This encourages the model to align the representations in the original script and the Latin script. We refer to this objective as transliteration language modeling (TLM) and the loss as $\mathcal{L}_{\text{TLM}}$.
The overall training objective combines the masked language modeling, sentence-level alignment, and token-level alignment objectives:
$\mathcal{L}=\mathcal{L}_{\text{MLM}}^{orig}+\mathcal{L}_{\text{MLM}}^{latn}+\mathcal{L}_{\text{SEQ}}+\mathcal{L}_{\text{TLM}}$
4 Experiments
4.1 General Setups
We use the Glot500 model (ImaniGooghari et al., 2023), a state-of-the-art multilingual encoder-only model pre-trained on more than 500 languages, as our source model for all our experiments. We fine-tune Glot500 on two groups of languages using the proposed transliteration-based post-training alignment method. The languages for each group are selected based on areal features so that they have some lexical overlap in different degrees and cover various scripts. We then evaluate the two resulting models on several downstream tasks in a zero-shot cross-lingual transfer manner. Apart from the standard transfer setting with English as the source language, we also evaluate the model’s transfer capabilities with three other source languages of different scripts for each language group.
Figure 2: Geographical distribution of languages selected in each group. Mediterranean-Amharic-Farsi are shown with circles, while South+East Asian Languages are shown with squares.
4.2 Languages, Data and Models
The two language groups are named as Mediterranean-Amharic-Farsi and South+East Asian Languages. We visualize each group’s geographical distribution of the selected languages in Figure 2. The languages within each group are spoken in adjacent areas and there is a long history of linguistic influence between them. For example, Arabic has had extensive contact with languages such as Turkish and Persian (Versteegh, 2001). The data for each language is sampled from the Glot500-c training dataset (ImaniGooghari et al., 2023). We sample 10% of the available data for each language, or a minimum of 10k sentences, whichever is larger. The data is then transliterated into Latin script using Uroman (Hermjakob et al., 2018). Table 1 shows each group’s languages and the number of sampled sentences. In total, Mediterranean-Amharic-Farsi consists of 10 languages, 5 scripts and around 16M sentences, while South+East Asian Languages consists of 10 languages, 7 scripts and around 4M sentences. We fine-tune Glot500 using our alignment method on each group separately. We then select the best checkpoint for each group by validating the checkpoints’ performance on the Tatoeba (Artetxe and Schwenk, 2019) sentence retrieval dataset, which contains 1000 English-aligned sentences. We compute the top-10 retrieval accuracy based on the cosine similarity of the averaged 8th-layer contextual embeddings. The best checkpoint for each group is regarded as our final aligned model.
Language Script Code Language Code Num. Sent. Mediterranean-Amharic-Farsi Macro Lang. Arabic Arab ara 2.4M Standard Arabic Arab arb 15k Moroccan Arabic Arab ary 10k Egyptian Arabic Arab arz 348k Macro Lang. Farsi Arab fas 1.8M Amharic Ethi amh 286k Greek Grek ell 2.2M Hebrew Hebr heb 1.8M Turkish Latn tur 2.9M Maltese Latn mlt 4M South+East Asian Languages Macro Lang. Chinese Hani zho 2.4M Classical Chinese Hani lzh 10k Yue Chinese Hani yue 48k Wu Chinese Hani wuu 22k Korean Hang kor 646k Lao Laoo lao 10k Lahu Latn lhu 10k Burmese Mymr mya 94k Tibetan Tibt bod 27k Thai Thai tha 773k
Table 1: Basic information and number of sampled sentences for each language in the two language groups.
SR-B Taxi1500 SIB200 NER POS
Latn Arab Grek Hebr Latn Arab Grek Hebr Latn Arab Grek Hebr Latn Arab Grek Hebr Latn Arab Grek Hebr
Glot500
tur_Latn 63.2 77.6 51.8 32.2 63.0 63.5 55.6 44.1 81.4 79.8 82.1 79.8 74.1 71.4 71.7 70.4 70.4 49.4 63.8 66.5
mlt_Latn 50.4 55.0 69.2 44.8 54.1 57.3 53.0 46.1 81.8 79.3 82.8 82.6 69.2 60.0 68.2 66.9 81.1 59.8 76.9 74.2
ell_Grek 48.6 58.6 src 40.8 64.0 61.4 src 44.1 77.7 74.4 src 81.7 72.7 72.1 src 72.6 86.1 58.3 src 67.8
heb_Hebr 21.8 27.8 33.8 src 35.4 44.9 39.0 src 77.6 73.9 81.7 src 48.9 58.7 52.6 src 68.3 70.1 60.6 src
amh_Ethi 52.8 64.6 51.2 33.2 7.2 10.4 12.7 15.1 73.1 74.9 74.1 74.8 43.8 52.5 54.0 46.6 66.5 65.1 64.3 73.6
ara_Arab - - - - - - - - - - - - 57.2 src 56.7 61.5 84.6 src 63.2 78.0
arz_Arab 24.8 33.6 52.8 44.8 35.1 43.1 45.2 42.7 79.7 src 81.8 80.4 58.4 75.1 63.8 65.2 - - - -
ary_Arab 15.2 16.4 29.0 28.4 35.8 40.3 41.6 39.6 79.9 80.2 84.0 82.0 - - - - - - - -
arb_Arab 14.6 23.0 29.0 32.2 - - - - 79.9 79.9 82.8 81.2 - - - - - - - -
fas_Arab 89.2 src 72.4 40.2 71.0 src 59.2 48.7 - - - - 49.7 66.2 58.2 50.6 71.5 67.2 60.9 72.0
Average 42.2 44.5 48.6 37.0 45.7 45.8 43.8 40.1 78.9 77.5 81.3 80.4 59.3 65.1 60.8 62.0 72.8 61.6 65.0 72.0
Ours
tur_Latn 81.0 91.2 77.8 49.4 64.8 65.1 54.6 38.2 85.6 86.0 84.8 85.6 77.0 73.1 76.9 73.2 73.6 52.8 66.4 68.0
mlt_Latn 85.6 93.4 90.4 59.4 66.6 60.9 58.4 42.5 86.2 85.8 84.6 85.4 75.2 72.0 73.9 75.2 83.1 63.0 77.1 77.8
ell_Grek 68.0 85.4 src 45.2 63.6 60.4 src 36.4 82.3 81.4 src 82.1 73.8 73.9 src 75.3 85.9 58.7 src 70.7
heb_Hebr 29.0 32.0 42.8 src 45.3 44.5 45.8 src 79.0 79.8 79.9 src 51.9 61.6 57.2 src 67.7 71.3 59.8 src
amh_Ethi 63.6 79.4 64.8 49.4 7.6 9.6 17.0 11.3 77.2 77.9 76.6 76.9 45.0 51.7 54.0 50.0 66.2 63.9 63.7 74.9
ara_Arab - - - - - - - - - - - - 59.9 src 60.4 65.1 65.3 src 62.3 77.7
arz_Arab 56.4 79.4 82.2 69.6 41.8 44.1 49.6 42.4 83.1 src 83.4 82.8 58.3 76.7 65.6 68.3 - - - -
ary_Arab 47.6 66.2 66.2 65.4 39.0 37.3 39.0 40.5 83.2 82.7 82.9 83.3 - - - - - - - -
arb_Arab 44.4 55.0 56.0 49.6 - - - - 82.8 83.3 83.1 83.1 - - - - - - - -
fas_Arab 89.6 src 87.0 57.8 71.9 src 63.2 37.5 - - - - 48.6 63.1 62.3 56.9 71.6 69.1 61.5 71.3
Average 62.8 72.7 70.9 55.7 50.1 46.0 46.8 35.5 82.4 82.4 82.2 82.7 61.2 67.4 64.3 66.3 73.3 63.1 65.1 73.4
Table 2: Cross-lingual transfer results across 5 tasks on the Mediterranean-Amharic-Farsi group. Columns represent the script of the source language (denoted with “src”), while rows represent the target languages. Results are averaged over 5 random seeds. For each source-target language pair, the best score is bolded. For each target language, we underline the best source transfer score for each task (for both Glot500 and our method).
4.3 Downstream Tasks
We evaluate the resulting aligned model for each group on several downstream tasks (described below). For each task, we use four different source languages, English and three other source languages belonging to the same group that use different scripts. The evaluation is done in a zero-shot cross-lingual transfer manner: we fine-tune the models on the train set of a given source language, select the best checkpoint, and compute the macro F1 score (except for SR-B where we compute the average top-10 retrieval accuracy) on the test sets of the remaining languages in each group. Note that no training step is needed for the retrieval task: we directly use the sentences from the source language as the queries and retrieve the most similar sentences in the target languages. For tasks that require additional fine-tuning, we report the results averaged over five different seeds. The downstream tasks are as follows (see details in §B):
SR-B
A sentence retrieval dataset where the parallel sentences are from the Bible. We compute the top-10 retrieval accuracy on 500 parallel sentences following ImaniGooghari et al. (2023).
Taxi1500
A multilingual text classification dataset covering more than 1500 languages with sentences from 6 topics (Ma et al., 2023).
SIB200
A multilingual text classification dataset covering more than 200 languages for 7 topics (Adelani et al., 2024).
NER
A multilingual sequence labeling dataset for named entity recognition  (Pan et al., 2017) that consists of articles annotated with 7 different tags, e.g., location, person, etc.
POS
A multilingual sequence labeling dataset for part-of-speech (POS) tagging (de Marneffe et al., 2021) consisting of sentences annotated with 17 universal POS tags, e.g., NOUN, ADJ, etc.
SR-B Taxi1500 SIB200 NER POS
Latn Hang Hani Thai Latn Hang Hani Thai Latn Hang Hani Thai Latn Hang Hani Thai Latn Hang Hani
Glot500
tha_Thai 45.4 47.2 42.6 src 64.1 63.8 73.6 src 82.0 83.1 83.1 src 4.3 2.8 7.5 src 55.0 29.8 49.0
kor_Hang 61.0 src 64.6 51.2 68.6 src 63.5 65.5 82.7 src 83.0 82.5 51.8 src 40.2 10.1 52.6 src 39.2
yue_Hani 24.0 31.6 65.8 44.4 64.0 62.8 68.0 62.1 84.7 84.4 src 86.9 24.0 37.4 69.1 16.4 38.8 49.3 78.5
wuu_Hani - - - - - - - - - - - - 35.1 58.3 62.4 16.1 - - -
zho_Hani 44.4 46.4 src 39.6 65.0 61.9 src 62.0 - - - - 23.6 32.7 src 15.0 40.1 49.9 src
lzh_Hani 63.4 65.8 75.8 43.2 57.3 60.5 61.5 56.1 - - - - 12.0 29.8 60.1 22.3 19.4 27.2 50.7
lao_Laoo 49.6 59.8 48.6 64.6 72.0 68.5 73.9 74.8 80.4 80.0 80.0 82.4 - - - - - - -
lhu_Latn 5.0 6.0 6.6 7.2 27.0 34.1 27.8 26.4 - - - - - - - - - - -
mya_Mymr 29.4 37.8 29.0 33.6 61.8 63.2 56.8 60.5 80.1 80.7 78.6 79.5 54.1 65.2 49.7 9.3 - - -
bod_Tibt 33.2 49.4 44.4 49.8 - - - - 70.0 68.8 65.1 72.7 36.5 42.8 50.0 25.5 - - -
Average 39.4 43.0 47.1 41.7 60.0 60.7 62.1 59.9 79.3 79.4 78.0 80.8 30.2 38.4 48.4 16.4 41.2 39.1 54.4
Ours
tha_Thai 45.2 85.6 55.2 src 66.3 66.7 67.8 src 86.6 85.9 82.9 src 3.5 2.6 8.6 src 50.6 30.3 48.6
kor_Hang 58.8 src 79.6 76.6 71.3 src 65.8 68.2 83.1 src 82.4 83.8 55.5 src 43.2 3.1 52.8 src 45.0
yue_Hani 71.4 91.8 98.8 89.8 64.5 66.4 69.2 67.9 87.2 84.7 src 89.0 21.0 36.1 70.1 15.8 29.3 38.6 79.3
wuu_Hani - - - - - - - - - - - - 45.5 56.7 64.7 1.7 - - -
zho_Hani 41.4 75.0 src 46.4 65.8 65.0 src 67.7 - - - - 21.5 35.9 src 15.6 32.2 44.1 src
lzh_Hani 37.4 48.6 67.6 37.0 63.4 58.3 61.4 55.2 - - - - 11.7 33.7 61.4 20.0 13.9 14.3 52.1
lao_Laoo 56.4 90.8 67.0 77.4 70.2 69.5 69.4 67.2 82.6 81.6 81.0 83.2 - - - - - - -
lhu_Latn 15.8 26.6 21.0 27.4 23.9 36.5 24.9 32.1 - - - - - - - - - - -
mya_Mymr 37.8 60.4 54.4 59.6 62.2 68.3 59.2 63.0 80.9 81.8 80.0 81.0 56.0 62.4 45.0 5.8 - - -
bod_Tibt 60.8 88.8 83.4 80.2 - - - - 70.5 73.5 71.4 71.4 40.3 43.1 41.6 1.1 - - -
Average 47.2 70.9 65.8 61.8 61.0 61.5 59.7 60.2 81.8 81.5 79.6 81.7 31.9 38.6 47.8 9.0 35.8 31.8 56.3
Table 3: Cross-lingual transfer results across 5 tasks on the South+East Asian Languages group. Columns represent the script of the source language (denoted with “src”), while rows represent the target languages. Results are averaged over 5 random seeds. For each source-target language pair, the best score is bolded. For each target language, we underline the best source transfer score for each task (for both Glot500 and our method)
5 Results and Analysis
We report the results of Glot500 and our post-trained aligned models on the downstream tasks in Table 2 for Mediterranean-Amharic-Farsi and in Table 3 for South+East Asian Languages. Overall, our aligned models outperform Glot500 across different tasks for both language groups, occasionally with a slight performance drop for certain source-target language combinations. In the following, we highlight our essential findings from the results.
Per-group performances differ slightly.
Starting with Mediterranean-Amharic-Farsi, we observe that the post-trained aligned model generally outperforms Glot500 on all downstream tasks. The SR-B task shows the most significant improvement, with the aligned model achieving, on average, more than 20% higher accuracy than Glot500 for all source languages. For other tasks, the aligned model also demonstrates a consistent, albeit smaller, improvement, with Glot500 occasionally outperforming the post-trained aligned model for specific source-target language pairs. However, we observe a more mixed performance for the South+East Asian Languages, especially for sequence labeling tasks, namely NER and POS, with the aligned models performing worse than Glot500 when transferring from more than half of the source languages. We hypothesize that this performance drop is primarily due to the transliteration process, which loses the semantic and contextual nuances and induces more token-level ambiguity for most of the languages (Amrhein and Sennrich, 2020; Liu et al., 2024a). This ambiguity makes token-level alignment more difficult. In contrast, the aligned model achieves consistent improvements in NER and POS for Mediterranean-Amharic-Farsi, where less token-level ambiguity is introduced as the languages are originally written in phonetic scripts, to which Latin also belongs.
SR-B Taxi1500 SIB200 NER POS
Latn Arab Grek Hebr Latn Arab Grek Hebr Latn Arab Grek Hebr Latn Arab Grek Hebr Latn Arab Grek Hebr
Glot500 42.2 44.5 48.6 37.0 45.7 45.8 43.8 40.1 78.9 77.5 81.3 80.4 59.3 65.1 60.8 62.0 72.8 61.6 65.0 72.0
MLM 50.2 53.2 55.4 40.2 47.7 43.8 46.9 31.8 82.6 82.4 82.5 80.4 61.7 68.5 63.3 66.5 71.9 62.4 63.7 72.0
MLM+SEQ 62.9 71.9 69.8 57.1 49.0 46.0 48.5 36.1 82.4 82.4 83.5 82.4 61.1 67.4 64.6 66.1 72.5 62.1 64.7 72.7
MLM+TLM 50.4 54.7 55.7 41.6 50.4 46.5 47.3 37.2 82.6 81.7 82.4 81.8 61.9 67.1 64.0 66.8 73.2 63.1 64.5 73.5
MLM+SEQ+TLM 62.8 72.7 70.9 55.7 50.1 46.0 46.8 35.5 82.4 82.4 82.2 82.7 61.2 67.4 64.3 66.3 73.3 63.1 65.1 73.4
Table 4: Ablation study results for Mediterranean-Amharic-Farsi. The columns represent the script of the source language. The results are averaged over all target languages. Bold (underlined): best (second-best) result.
SR-B Taxi1500 SIB200 NER POS
Latn Hang Hani Thai Latn Hang Hani Thai Latn Hang Hani Thai Latn Hang Hani Thai Latn Hang Hani
Glot500 39.4 43.0 47.1 41.7 60.0 59.2 60.7 58.2 79.3 79.4 78.0 80.8 30.2 38.4 48.4 16.4 41.2 39.1 54.4
MLM 45.8 54.3 54.7 51.4 58.5 60.2 60.4 60.3 80.9 80.3 79.4 81.6 28.7 37.9 46.7 7.2 29.5 27.3 54.2
MLM+SEQ 47.0 63.8 61.0 56.1 60.6 59.9 60.0 61.0 81.8 79.3 79.3 80.3 30.9 39.3 46.5 7.7 31.4 30.9 55.4
MLM+TLM 45.4 55.7 57.8 54.5 61.4 60.6 59.5 60.0 81.9 81.5 79.5 81.5 31.0 39.0 47.6 9.1 34.6 32.1 54.8
MLM+SEQ+TLM 47.2 70.9 65.8 61.8 61.0 61.5 59.7 60.2 81.8 81.5 79.6 81.7 31.9 38.6 47.8 9.0 35.8 31.8 56.3
Table 5: Ablation study results for South+East Asian Languages. The columns represent the script of the source language. The results are averaged over all target languages. Bold (underlined): best (second-best) result.
Source languages matter.
We observe that the performance can vary significantly for both groups of languages depending on the source language used for transfer. This phenomenon occurs due to the script and language similarity between specific source and target languages. In general, transferring from in-group high-resource languages performs better than transferring from English. Taking the SR-B task for example, for Mediterranean-Amharic-Farsi, the best performance is achieved when transferring from Farsi. For the South+East Asian Languages, the best performance is achieved when transferring from Korean. The text classification tasks generally show less variation in performance, though transferring from Hebrew achieves the worst performance for Mediterranean-Amharic-Farsi in Taxi1500. For the NER and POS tasks, transferring from Arabic and Hebrew achieves the best performance for Mediterranean-Amharic-Farsi while transferring from Chinese achieves the best performance for both tasks in South+East Asian Languages. Nevertheless, comparing our aligned models against Glot500, the performance generally improves for most source languages. This indicates that our proposed transliteration-based post-training method effectively improves the alignment between related languages and further boosts performance when a proper source language is chosen.
5.1 Ablation Study
We perform an ablation study to investigate the impact of different training objectives on the performance of the post-trained aligned models. Starting from the base Glot500 model, we apply different combinations of the training objectives: masked language modeling (MLM), sentence-level alignment (SEQ), and token-level alignment (TLM). There are four different combinations in total (MLM+SEQ is the training objective of TransliCo (Liu et al., 2024a)). Note that we do not consider the variant where the MLM objective is missing since MLM is important to preserve the language modeling capability. We report the average performance of using four different source languages across all target languages for each language group in Table 4 and Table 5.
The lone MLM objective already provides some slight improvement over the base Glot500 model. We hypothesize this is due to the benefit of specializing the model to a small group of languages. When the sequence-level alignment (SEQ) objective is included, the performance is generally further improved compared with the MLM variant, especially for the retrieval task. This is not surprising as the SEQ improves the sequence-level alignment. When instead the token-level alignment (TLM) objective is included, there is a slight improvement in the performance for most tasks compared to the MLM+SEQ objective, except for the retrieval task. We also observe that text classification tasks show the least variation in performance for most source languages. This is especially the case for SIB200, which seems to be the easiest task for the models.
Our whole objective (MLM+SEQ+TLM) generally performs better than the other models by combining sequence and token-level alignment benefits. For Mediterranean-Amharic-Farsi, though MLM+SEQ+TLM performs on par with the MLM+SEQ objective on NER, it achieves better performance on POS. Similarly, our complete training objective outperforms MLM+SEQ on both NER and POS for South+East Asian Languages. When comparing the performance of the full training objective against MLM+TLM, we observe an equal performance across different tasks, except for the retrieval task, where the full training objective noticeably outperforms MLM+TLM. Even though SEQ objective is the most critical for improving the retrieval task, we also observe a performance increase in the retrieval task when TLM is included. This indicates that our post-training alignment method, with TLM, effectively improves both the sequence and token-level alignment.
6 Conclusion
In this work, we propose a transliteration-based post-training method that contains both sequence and token-level objectives to improve the cross-lingual/script alignment of mPLMs and thus boost their zero-shot cross-lingual transfer performance. We apply our post-training method to fine-tune Glo500 on two language groups that share areal features and have extensive lexical overlap. Our extensive experiments using different source languages show that our aligned models consistently outperform the original Glot500 model. In particular, our method enhances the alignment between related languages and, therefore, improves cross-lingual transfer between these languages. We also analyze the impact of different training objectives and show that the sequence and token-level alignment objectives are both critical for achieving the best performance across different tasks.
Limitations
Even though the mPLM is fine-tuned with our method, where the transliterated text is used as an auxiliary input, the mPLM has only seen the Latin transliterations during its pre-training phase. This can limit the performance of the post-trained aligned models, especially for languages with complex scripts. An extension of this work could expand the vocabulary to include the subwords from Latin transliterations as done by Liu et al. (2024b) before fine-tuning or continued-pre-train the model on the transliterated text so that the models can be more effective in modeling the transliterated data.
We are further limited by the transliteration process, which only partially captures the phonetic and semantic information of the original text, especially for languages with significantly different scripts from the Latin script. This leads to a loss of information during the alignment process, which can negatively impact the performance of the post-trained aligned models. Future work could improve the transliteration process to better capture the linguistic properties of the original text.
Acknowledgments
This research was supported by DFG (grant SCHU 2246/14-1) and The European Research Council (NonSequeToR, grant #740516). We would like to thank Leibniz Supercomputing Center (LRZ) and Munich Center for Machine Learning (MCML) for providing the computational resources.
Appendix A Training Details
For the MLM objective, we use the normal masking probability of 15%. We use the AdamW optimizer (Kingma and Ba, 2015; Loshchilov and Hutter, 2019) with an initial learning rate of $82e-5,\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=1e-$ 8, weight decay of $0.01$, and a linear learning rate scheduler with no warm-up steps. We use FP16 mixed precision training (Micikevicius et al., 2018). Each batch contains sentence pairs in the original and Latin scripts, with a maximum sequence length of 512 tokens. We set the per-device batch size to 16, gradient accumulation steps to 8, and use four NVIDIA A100 80GB GPUs. This leads to an effective batch size of 512. The Mediterranean-Amharic-Farsi model is trained for 2 epochs, while the South+East Asian Languages model is trained for 8 epochs. We use the HuggingFace Transformers library (Wolf et al., 2020) for all experiments. The trainings take around 30 hours for both groups. We store the model checkpoints every 2000 steps.
For the sequence-level contrastive learning objective, we unify all the per-device batch sentence embeddings into a global batch in order to have a larger amount of negative samples. For the token-level alignment objective, differently from the original TLM, we do not reset the positional embeddings for the second sentence in a given pair.
Appendix B Downstream Tasks Fine-Tuning
For the downstream tasks that require fine-tuning, we further fine-tune the post-trained aligned models on the training set of a given source language, select the best checkpoint with early stopping based on the f1 score on the source language’s validation set, and evaluate the macro F1 score on the test sets of the remaining languages in each group. The fine-tuning results are averaged over five different seeds. All models are fine-tuned on a single GPU. Unless otherwise stated, the same optimizer and scheduler settings are used as in the post-training alignment phase.
SR-B
We use the same setup as ImaniGooghari et al. (2023), by calculating the top-10 retrieval accuracy on 500 parallel sentences from the Bible. This task does not require any fine-tuning, and the retrieval is done by computing the cosine similarity between the average 8th-layer contextual embeddings of the source and target sentences.
Taxi1500
A multilingual text classification dataset covering more than 1500 languages formed by classifying 1077 bible verses into six topics (Ma et al., 2023). The learning rate is set to $51e-$ 5, and we fine-tune the models for 40 epochs with a batch size of 32 and a maximum sequence length of 100 tokens.
SIB200
A multilingual text classification dataset covering more than 200 languages formed by classifying 1004 article sentences into 7 topics (Adelani et al., 2024). We use the same fine-tuning setup as the Taxi1500 task, except for the maximum sequence length, which is set to 160 tokens.
NER
We evaluate the models on the WikiAnn named entity recognition dataset (Pan et al., 2017), a multilingual dataset consisting of articles annotated with 7 different tags. We set the learning rate to $52e-$ 5, the batch size to 32, gradient accumulation steps to 2, and the maximum sequence length to 256 tokens. We fine-tune the models for 5 epochs.
POS
We evaluate the models on the Universal Dependencies (UD) v2.11 part-of-speech tagging dataset (de Marneffe et al., 2021), a multilingual dataset consisting of sentences annotated with 17 universal POS tags. We use the same fine-tuning setup as the NER task, except for number of epochs, which is set to 10.
Appendix C Vocabulary Analysis
We compare the coverage of the Glot500 vocabulary in the original and transliterated corpora by tokenizing the fine-tuning datasets and counting the number of unique tokens. The results are shown in Table 6. As expected, the transliterated text is represented by a smaller part of the vocabulary leading to text being broken down into smaller subwords. This result suggests that the performance of the post-trained aligned models could be further improved by extending the vocabulary of the pre-trained model based on the transliterated corpus.
Language Group Original Tokens Transliterated Tokens Mediterranean-Amharic-Farsi 209K 125K South+East Asian Languages 120K 88K
Table 6: Number of unique tokens in the Glot500 vocabulary covered in the original and transliterated corpora for each language group. A smaller set of unique tokens is used after transliterating the corpora into the common Latin script.
Appendix D Full Ablation Results
We provide the full results of the other models trained with different combinations of training objectives for both language groups in Table 7, Table 8, Table 9, Table 10, Table 11, and Table 12.
SR-B Taxi1500 SIB200 NER POS
Latn Arab Grek Hebr Latn Arab Grek Hebr Latn Arab Grek Hebr Latn Arab Grek Hebr Latn Arab Grek Hebr
tur_Latn 78.6 89.4 67.2 38.6 65.1 65.1 58.9 36.2 86.8 86.2 84.8 82.9 77.2 73.1 76.5 73.9 73.2 51.1 66.5 68.7
mlt_Latn 76.2 78.6 80.2 47.2 61.3 56.6 59.5 35.6 84.1 85.8 84.8 83.0 75.2 73.3 71.6 74.7 81.6 62.6 77.5 75.8
ell_Grek 57.2 64.2 src 36.2 61.5 60.5 src 32.4 82.9 81.3 src 79.9 73.8 74.1 src 75.0 85.3 57.0 src 63.0
heb_Hebr 25.6 28.0 32.2 src 43.8 43.3 45.8 src 78.6 77.2 79.2 src 50.8 60.5 55.9 src 65.8 70.8 57.7 src
amh_Ethi 62.8 77.8 60.8 38.2 4.8 5.1 12.5 11.6 77.8 79.4 76.6 76.0 43.8 53.2 52.7 49.8 64.5 64.4 60.5 74.7
ara_Arab - - - - - - - - - - - - 58.2 src 59.6 66.6 62.9 src 59.8 77.2
arz_Arab 30.4 42.4 59.4 51.2 39.5 38.4 47.8 37.3 84.4 src 82.9 80.3 58.9 77.6 65.3 67.4 - - - -
ary_Arab 19.4 23.4 34.0 36.0 35.2 37.8 39.2 36.0 82.9 81.7 85.3 81.5 - - - - - - - -
arb_Arab 12.6 22.4 30.8 33.4 - - - - 83.0 85.0 83.4 79.2 - - - - - - - -
fas_Arab 89.8 src 79.2 40.8 70.1 src 64.6 33.8 - - - - 55.6 67.8 61.7 58.0 69.6 68.6 60.2 72.5
Average 50.2 53.2 55.4 40.2 47.7 43.8 46.9 31.8 82.6 82.4 82.5 80.4 61.7 68.5 63.3 66.5 71.9 62.4 63.7 72.0
Table 7: Results for the Mediterranean-Amharic-Farsi model with the MLM objective. Columns represent the script of the transfer source language (denoted with “src”), while rows represent the target languages. Results are averaged over 5 random seeds. For each target language, we underline the best source transfer score for each task.
SR-B Taxi1500 SIB200 NER POS
Latn Arab Grek Hebr Latn Arab Grek Hebr Latn Arab Grek Hebr Latn Arab Grek Hebr Latn Arab Grek Hebr
tur_Latn 80.6 90.0 77.2 52.4 68.1 70.5 62.2 40.0 86.8 87.4 86.8 85.9 78.0 73.0 77.2 72.7 73.5 50.0 66.7 68.2
mlt_Latn 83.2 90.2 87.8 62.8 64.1 59.2 59.2 40.3 84.8 85.2 85.4 84.3 74.5 71.3 72.1 75.0 82.4 61.9 76.8 75.9
ell_Grek 67.0 83.0 src 47.4 61.6 61.8 src 38.9 83.1 82.0 src 82.2 73.1 73.6 src 75.1 85.9 57.5 src 68.6
heb_Hebr 32.0 37.4 45.4 src 45.9 46.0 45.9 src 79.8 77.5 80.1 src 50.4 60.6 57.0 src 66.6 71.1 58.8 src
amh_Ethi 57.4 71.0 57.8 46.6 4.8 6.2 14.6 13.7 76.4 79.2 78.8 77.5 41.9 52.3 54.1 49.5 65.1 63.6 63.1 74.1
ara_Arab - - - - - - - - - - - - 56.5 src 61.9 65.8 63.9 src 61.7 78.0
arz_Arab 59.0 79.2 82.6 71.4 39.1 40.3 49.3 43.0 83.6 src 83.9 81.8 59.1 75.6 66.9 67.3 - - - -
ary_Arab 52.6 68.4 68.4 68.0 36.9 38.2 39.5 38.8 82.3 81.5 84.8 82.8 - - - - - - - -
arb_Arab 47.6 56.0 56.2 52.2 - - - - 82.6 84.0 85.0 82.6 - - - - - - - -
fas_Arab 87.0 src 83.0 56.2 71.2 src 68.6 38.0 - - - - 55.1 65.2 63.0 57.4 69.9 68.2 60.7 71.3
Average 62.9 71.9 69.8 57.1 49.0 46.0 48.5 36.1 82.4 82.4 83.5 82.4 61.1 67.4 64.6 66.1 72.5 62.1 64.7 72.7
Table 8: Results for the Mediterranean-Amharic-Farsi model with the MLM+SEQ objective. Columns represent the script of the transfer source language (denoted with “src”), while rows represent the target languages. Results are averaged over 5 random seeds. For each target language, we underline the best source transfer score for each task.
SR-B Taxi1500 SIB200 NER POS
Latn Arab Grek Hebr Latn Arab Grek Hebr Latn Arab Grek Hebr Latn Arab Grek Hebr Latn Arab Grek Hebr
tur_Latn 78.2 89.4 66.2 40.0 64.0 66.2 59.9 41.2 86.8 86.2 84.9 85.2 76.7 72.6 76.4 73.3 73.3 52.4 66.9 68.9
mlt_Latn 76.8 82.0 82.2 48.0 66.8 61.5 60.2 45.0 86.3 85.1 85.7 83.6 75.8 72.0 72.6 75.6 83.1 64.0 77.3 78.7
ell_Grek 56.6 64.8 src 34.0 65.1 66.5 src 40.0 82.5 81.5 src 81.8 73.6 73.7 src 75.5 85.8 58.3 src 67.4
heb_Hebr 26.2 27.8 33.2 src 47.1 43.9 41.4 src 77.6 77.6 80.7 src 51.9 60.4 56.6 src 67.5 70.8 59.2 src
amh_Ethi 60.0 78.2 57.0 37.4 7.5 8.5 17.0 10.6 77.1 77.6 76.7 76.4 46.2 51.7 54.0 46.6 65.7 63.4 62.5 75.3
ara_Arab - - - - - - - - - - - - 60.1 src 59.3 66.8 65.7 src 60.0 77.9
arz_Arab 30.2 45.4 61.2 54.4 42.8 40.9 49.4 42.1 83.9 src 82.2 82.6 57.6 76.0 66.2 68.9 - - - -
ary_Arab 21.8 27.0 35.6 39.6 38.3 37.8 39.6 39.0 83.6 81.1 82.5 81.5 - - - - - - - -
arb_Arab 13.8 23.6 31.2 36.2 - - - - 83.1 82.9 84.0 81.6 - - - - - - - -
fas_Arab 90.0 src 79.6 43.4 72.0 src 63.5 42.2 - - - - 53.5 63.2 63.0 60.7 71.6 69.6 60.9 73.0
Average 50.4 54.7 55.7 41.6 50.4 46.5 47.3 37.2 82.6 81.7 82.4 81.8 61.9 67.1 64.0 66.8 73.2 63.1 64.5 73.5
Table 9: Results for the Mediterranean-Amharic-Farsi model with the MLM+TLM objective. Columns represent the script of the transfer source language (denoted with “src”), while rows represent the target languages. Results are averaged over 5 random seeds. For each target language, we underline the best source transfer score for each task.
SR-B Taxi1500 SIB200 NER POS
Latn Hang Hani Thai Latn Hang Hani Thai Latn Hang Hani Thai Latn Hang Hani Thai Latn Hang Hani
tha_Thai 39.6 47.8 46.4 src 65.2 65.1 70.6 src 85.5 84.3 83.8 src 1.9 0.4 8.7 src 47.1 29.2 48.2
kor_Hang 74.4 src 70.0 65.4 63.3 src 62.0 65.4 82.8 src 83.0 83.8 56.2 src 40.7 2.9 52.9 src 38.2
yue_Hani 38.2 55.8 82.4 67.6 68.6 66.4 70.2 69.4 86.0 85.5 src 88.3 16.8 33.7 70.4 11.8 20.3 30.7 78.6
wuu_Hani - - - - - - - - - - - - 32.2 54.2 62.8 1.3 - - -
zho_Hani 46.0 49.8 src 44.4 64.8 67.8 src 67.8 - - - - 17.7 33.1 src 11.0 20.6 33.9 src
lzh_Hani 62.2 70.4 74.8 44.6 60.0 58.5 64.4 57.6 - - - - 8.2 33.3 62.0 14.7 6.7 15.2 51.6
lao_Laoo 53.6 78.4 53.8 72.2 67.9 67.4 72.6 69.3 83.0 80.8 81.2 83.9 - - - - - - -
lhu_Latn 6.4 8.8 7.6 10.4 20.2 31.1 21.7 29.6 - - - - - - - - - - -
mya_Mymr 32.8 43.2 32.8 43.2 57.8 65.2 61.3 62.9 79.6 79.6 80.5 79.6 56.6 68.6 38.7 3.6 - - -
bod_Tibt 59.8 80.4 70.2 64.0 - - - - 68.6 71.4 68.5 72.1 39.7 42.0 43.2 5.5 - - -
Average 45.8 54.3 54.7 51.4 58.5 60.2 60.4 60.3 80.9 80.3 79.4 81.6 28.7 37.9 46.7 7.2 29.5 27.3 54.2
Table 10: Results for the South+East Asian Languages model with the MLM objective. Columns represent the script of the transfer source language (denoted with “src”), while rows represent the target languages. Results are averaged over 5 random seeds. For each target language, we underline the best source transfer score for each task.
SR-B Taxi1500 SIB200 NER POS
Latn Hang Hani Thai Latn Hang Hani Thai Latn Hang Hani Thai Latn Hang Hani Thai Latn Hang Hani
tha_Thai 49.0 75.6 55.6 src 66.4 67.1 71.2 src 86.3 83.6 84.3 src 2.7 1.7 8.3 src 47.8 27.6 48.3
kor_Hang 58.6 src 72.2 70.8 67.3 src 61.6 68.3 82.9 src 83.3 83.4 55.8 src 41.2 3.6 53.1 src 41.8
yue_Hani 60.2 84.4 97.4 81.0 66.2 61.9 67.7 67.2 86.9 83.7 src 88.0 19.3 35.0 70.5 12.2 23.0 37.9 79.4
wuu_Hani - - - - - - - - - - - - 44.2 58.6 63.9 2.0 - - -
zho_Hani 45.2 67.4 src 44.4 66.2 64.5 src 68.0 - - - - 19.9 36.2 src 13.2 23.5 41.2 src
lzh_Hani 36.6 37.2 55.8 30.0 65.0 59.8 60.1 58.4 - - - - 9.3 31.4 62.4 18.8 9.6 16.8 52.1
lao_Laoo 55.2 81.6 60.2 72.6 70.9 66.2 72.9 67.9 82.8 80.9 82.2 83.5 - - - - - - -
lhu_Latn 19.4 23.4 18.0 22.4 23.1 36.7 27.6 30.7 - - - - - - - - - - -
mya_Mymr 40.2 58.0 49.0 50.6 59.5 63.4 58.5 66.3 81.8 79.6 78.0 78.7 57.2 65.6 43.2 2.4 - - -
bod_Tibt 58.6 83.4 79.8 77.4 - - - - 70.2 68.9 68.9 68.1 38.9 46.2 36.3 1.7 - - -
Average 47.0 63.8 61.0 56.1 60.6 59.9 60.0 61.0 81.8 79.3 79.3 80.3 30.9 39.3 46.5 7.7 31.4 30.9 55.4
Table 11: Results for the South+East Asian Languages model with the MLM+SEQ objective. Columns represent the script of the transfer source language (denoted with “src”), while rows represent the target languages. Results are averaged over 5 random seeds. For each target language, we underline the best source transfer score for each task.
SR-B Taxi1500 SIB200 NER POS
Latn Hang Hani Thai Latn Hang Hani Thai Latn Hang Hani Thai Latn Hang Hani Thai Latn Hang Hani
tha_Thai 37.6 46.8 45.6 src 64.6 63.7 69.4 src 86.5 84.5 85.0 src 1.7 2.2 9.4 src 50.4 30.6 47.4
kor_Hang 72.8 src 75.0 67.6 66.9 src 61.4 66.0 83.6 src 82.8 83.7 56.0 src 43.4 4.0 52.8 src 41.2
yue_Hani 43.0 67.2 88.8 76.6 69.1 62.7 66.3 65.2 87.1 86.0 src 89.1 19.4 34.7 70.6 13.4 27.4 38.9 79.4
wuu_Hani - - - - - - - - - - - - 44.5 56.6 63.0 3.6 - - -
zho_Hani 45.0 49.4 src 44.0 68.3 63.2 src 69.2 - - - - 19.8 34.0 src 12.4 30.3 43.6 src
lzh_Hani 57.6 69.6 80.0 48.4 64.0 59.7 61.6 55.4 - - - - 12.1 35.4 61.1 17.2 12.1 15.3 51.3
lao_Laoo 57.2 77.4 54.2 71.8 72.2 68.9 72.6 70.4 81.9 82.5 81.4 83.6 - - - - - - -
lhu_Latn 8.0 10.4 9.4 13.2 22.6 38.4 25.0 31.9 - - - - - - - - - - -
mya_Mymr 33.0 44.2 39.4 47.4 63.6 67.9 60.1 61.6 80.9 81.2 80.0 80.3 54.3 66.3 40.5 5.4 - - -
bod_Tibt 54.6 80.6 70.0 67.2 - - - - 71.2 73.5 68.5 70.7 40.4 44.2 44.9 8.0 - - -
Average 45.4 55.7 57.8 54.5 61.4 60.6 59.5 60.0 81.9 81.5 79.5 81.5 31.0 39.0 47.6 9.1 34.6 32.1 54.8
Table 12: Results for the South+East Asian Languages model with the MLM+TLM objective. Columns represent the script of the transfer source language (denoted with “src”), while rows represent the target languages. Results are averaged over 5 random seeds. For each target language, we underline the best source transfer score for each task.